In the vertices of a square of side 2x10-9 m four protons are placed. A fifth proton is initially on the perpendicular to the square by its center, at a distance of 2x10-9 m from it. Calculate (a) the minimum inicial velocity that the fifth proton needs to reach the center of the square, (b) its inicial and final accelerations. (c) Make a graph of proton potencial energy in function of its distance to the center of the square.

Please answer step by step and explain each step so I can understand.
Answer #1

flood 1 F Estre xlolm = 20 2x109 m=20A a proton 5th experiencing electric potenteal Energy by formalas- Call sides have beenKos Imor 4x1007 254 22 I 2X10 anto (1062 1056 1 mp = 1.67X10-27 kg x/0-27 $X 167 Å 20 2 - (1.6870-1972 Tox 10-704 *fit 10x/0-12 Sog mai - 4K 22 2 22 ū (r = 1056 Å 1.67820² + ai = 4x9x10°x (1.6x109972 (1056x1070)2 55 T 1067X1627 i = 0.0886*10-9 - I ai